The 5th term of a geometric Sequence is 8 and the 9th term Is 16128.find the first term


1
Expert's answer
2021-10-01T13:10:47-0400

"First \ term \ GP = a\\
Second\ term = ar\\
Third \ term = ar\\^2\\
5th\ term = ar\\^4\\
9th\ term = ar\\^8\\
ar\\^4 = 8\\
ar\\^8 = 16128
\\
a = \frac{16128}{r\\^8}\\
\frac{16128}{r\\^8} \times r\\^4 = 8"




"\frac{16128}{r\\^4} = 8 \\ \\ 8r\\^4 = 16128 \\ \frac{16128}{8} = r\\^4
\\ 2016 = r\\^4
\\ r = 6.70 \\a = \frac{8}{r\\^4} \\
 a = \frac{8}{(6.7)\\^4}
\\ a = 0.00397"